Customers using a self-service soda dispenser take an average of 12 ounces of soda with an SD of 4 ounces. Assume that the amount would be normally distributed.

a. What is the chance that a randomly selected customer takes over 10 ounces of soda?
b. What is the chance that a randomly selected customer takes between 13 and 14 ounces of soda?
c. What is the chance that the next 100 customers will take an average of less than 12.24 ounces?
1
Expert's answer
2019-10-23T08:35:58-0400

Part a). We asked to calculate "P(X > 10)" . Let's change the variable


"Z = \frac{{X - \mu }}{\sigma } = \frac{{X - 12}}{4}"

Then "Z" will be distributed according to standard normal distribution "Z \sim N(0,1)". Let's rewrite "P(X > 10)" as


"P(X > 10) = P(Z > \frac{{10 - 12}}{4}) = P(Z > - \frac{1}{2}) = P(Z < \frac{1}{2})"

(at the last one we use simmetry of the standard normal distribution about "x = 0"). Now let's use the following table




and get "P(Z < \frac{1}{2}) \approx 0.6915", thus

"P(X > 10) \approx 0.6915"

Part b). We asked to calculate "P(13 < X < 14)". Using the same variable "Z" we can rewrite it as


"P(13 < X < 14) = P(\frac{1}{4} < Z < \frac{1}{2}) = P(Z < \frac{1}{2}) - P(Z < \frac{1}{4})"

Using the table "P(Z < \frac{1}{2}) \approx 0.6915" and "P(Z < \frac{1}{4}) \approx 0.5987" . Thus


"P(13 < X < 14) \approx 0.6915 - 0.5987 \approx 0.0928"

Part c). Let "{X_i}" be the amount of ounces of soda that the i-th customer will get, and "{X_i} \sim N(12,4)" . Let's use the following lemma: Let "\xi \sim N({\mu _1},{\sigma _1})" and "\eta \sim N({\mu _2},{\sigma _2})" be the two normally distributed random values. Then "(\xi + \eta ) \sim N({\mu _1} + {\mu _2},\sqrt {\sigma _1^2 + \sigma _2^2} )" . You may find two different proofs of this here

https://en.wikipedia.org/wiki/Sum_of_normally_distributed_random_variables

Using this lemma we see that


"Y = \sum\limits_{k = 1}^{100} {{X_k}} \sim N(1200,40)"

Now we need to find "P(\frac{Y}{{100}} < 12.24) = P(Y < 1224)" . Let's change the variable


"Z = \frac{{Y - 1200}}{{40}}"

and we get


"P(\frac{Y}{{100}} < 12.24) = P(Z < \frac{3}{5})"

Using the table "P(Z < \frac{3}{5}) \approx 0.7257" thus


"P(\frac{Y}{{100}} < 12.24) \approx 0.7257"

This is the answer.